OpenStudy (anonymous):

a professional lawn care service can mow 2 3/4 acres of lawn in one hour. how many hours would it take them to mow a lawn that is 6 /78 acres?

OpenStudy (anonymous):

2 3/4 acres = 11/4 acres in 1 hour. To mow 1 acre it takes 4/11 hours 6/78 = 1/12 acres 1/12 of the lawn will be mowed in 4/11/12 = 4/(11*12) = 1/33 hours

OpenStudy (anonymous):

oh i messed up on the problem its 6 7/8. im sorry

OpenStudy (anonymous):

2 3/4 acres = 11/4 acres in 1 hour. To mow 1 acre it takes 4/11 hours 6 7/8 is 55/8 For one acre it takes 4/11 hours For 55/8 we will 4/11 * 55/8 = 5/2 = 2.5 hours

OpenStudy (anonymous):

thanks rita i seen where i made my mistake. thanks!